OpenStudy (anonymous):

Solve for x e^(8x+5)=7 Give your answer to 4 decimal places.

OpenStudy (anonymous):

ıntroduce ln

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[e^x=b\] \[x \ln e=\ln b\]

OpenStudy (anonymous):

\[(8x+5)\ln e=\ln 7\]

OpenStudy (anonymous):

I never got it right. Thanks anyways!